Circular No. 5955

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

4U 1145–619 R. B. Wilson, M. H. Finger, M. Stollberg, T. A. Prince, J. M. Gruns-

feld, B. A. Harmon, and S. N. Zhang report for the Compton Gamma Ray Observatory BATSE team: “An outburst from an x-ray pulsar is currently being observed. Pulsations were first detected on Mar. 12.5 UT. The measured barycentric pulse period was 293.4464 ± 0.0016 s on Mar. 19.0. The location, determined on Mar. 19 using earth-occultation mapping, is α = 11^{h}48^{m}.6, δ = −62^{o}03′ (equinox 2000.0), with an error radius of 0^{o}.2. This location is consistent with either of the transient x-ray binary systems 4U 1145–619 or 1E 1145.1–6141. Based on previously reported periods of 292 and 297 s, respectively, the source is most likely 4U 1145–619. The period rate of change measured between Mar. 17 and 21 was (−2.43 ± 0.34) × 10^{−7}. The pulse shape averaged over Mar. 16–20 consists of a narrow peak with a phase width of ∼ 0.35, immediately followed by a deep ‘well’ with a width of ∼ 0.1 in phase. The total flux in the band 20–40 keV on Mar. 19 was 0.5 Crab. The phase-averaged pulsed flux on that day was 0.3 Crab (total). These flux values each have uncertainties of ∼ 10 percent. The source intensity has remained within 10 percent of these values through Mar. 21.0.”

GRS 1009–45 J. Orosz, S. Barnes, and C. Bailyn, Astronomy Department, Yale Uni-

versity, report: “We have observed the x-ray nova GRS 1009–45 (cf. IAUC 5864, 5890) over a period of 18 nights in February using the Cerro Tololo Interamerican Observatory 0.9-m telescope (+ CCD imager). Point spread function fitting was used to separate the nova from an 18th-mag star located 1′′.5 to its southeast. On Feb. 7.12 UT, the nova was at V = 16.15 ± 0.02 and B − V = +0.23. Thereafter, the source brightness decreased at an increasing rate until Feb. 21.62, when its magnitude was V = 18.94 ± 0.02. The next observation at Feb. 22.18, 163 days after the initial outburst, showed V = 18.46, and subsequently the source continued to brighten at a rate of ∼ 0.02 mag/hr; by our last observation on Feb. 24.34, it had attained V = 17.43 ± 0.02. This secondary outburst is reminiscent of other x-ray novae, for example GRO J0422+32 (IAUC 5842, 5901, 5923). Further monitoring of this source is encouraged.”

1994 March 23 Daniel W. E. Green